If range is 2 times of height calculate horizontal rangr in term of uand g

Grade:11

1 Answers

Vikas TU
7 years ago
If range is 2 times of height then,
R = 2*H
that is to say that,
2usinthetha/g = 2*(u^2sin^thetha/g).
1 = usinthetha.................(1)
Range = 2usinthetha/g.
From eqn. (1)
Range = 2/g = 0.2 meter.